### v1

# Fully quantum stochastic entropy production

Abstract
Building on the approach of stochastic thermodynamics, we define entropy production for arbitrary quantum processes, using Bayesian tools to capture the inherent subjective element. First, we show that the classical expression for average entropy production involves only comparisons of statistics at the input or output, and thus can be formally generalized by replacing probability distributions by quantum states. To go beyond the mere average, we construct an entropy production operator, that generalizes the value of entropy to the non-commutative case. The operator is Hermitian; its average is non-negative, so that the Second Law takes its usual form $\left\langle\Sigma\right\rangle\geqslant 0$; it formally satisfies the Jarzynski equality; and recovers the classical expression for classical processes. We also present features whose fully quantum version is not a straightforward translation of the classical one: these may point to the core difference between classical and quantum thermodynamics.
I Introduction
Entropy is one of the most profound concepts in science. First introduced as a precise concept in thermodynamics by Clausius in the mid-19th century, it was destined to play a crucial role in many other areas of science. After reappearing in Shannon’s information theory in the 1940s, it has since found surprising applications in diverse fields ranging from cosmology to biology and economics. The last decades have seen several attempts of unification, starting with Landauer’s and Bennett’s analysis of the Maxwell’s demon.
Defining entropy for a given physical process is more than a matter of calculation: it requires a dose of expert wisdom. Reviews are still being written to guide the process [1]. The debate is even more open in the quantum realm [2]. Strasberg and Schindler considered no less than eleven well-accepted definitions of entropy in thermodynamics: they showed that they are not equivalent, and that some of them violate the Second Law in some cases [3]. In a parallel line of work, generalizations from the formalism of classical stochastic thermodynamics have led to consider the possibility that entropy takes complex values [4], or even that its average may be complex [5, 6], thus requiring a formulation of the Second Law different from the usual statement that the production of entropy must be non-negative. This work aims to address these challenges by formulating a fully quantum notion of entropy production that takes real values and has non-negative average, avoiding radical departures from the conventional operational interpretation.
Entropy inherently reflects a degree of subjectivity, as it depends on the knowledge and control available to the observer. This also implies that it is not straightforward to discriminate different definitions by way of experiment: since there is no universal entropy-meter, everyone can in principle design a device that measures their favorite quantity, especially for processes that are not amenable to energetics considerations. Notwithstanding, we argue that this situation does not justify the coexistence of multiple, incommensurate definitions of entropy. Instead, we should aim to converge on a single, unified definition—one in which the role of knowledge is made explicit.
This paper contributes to this research agenda by proposing a framework for defining the entropy production of arbitrary quantum processes. Drawing inspiration from stochastic thermodynamics, we conceptualize entropy production as a statistical comparison between a physical process and its associated reverse process. The subjective aspect arises in defining the reverse process via Bayesian retrodiction, which requires a prior. Our approach enforces that the Second Law holds in its standard form and ensures consistency with the classical formula for classical processes.
After reviewing the classical formalism (Section II), we present two definitions. For the first definition (Section III), we generalize the classical formula for average entropy production from Bayesian retrodiction, replacing probability distributions with quantum states. This approach recovers many specific definitions proposed over the years as special cases. However, it leaves open the question of what exactly is being averaged.
For the second definition (Section IV), we construct a Hermitian entropy production operator using suitable states-over-time to describe both the physical process and its Bayesian reverse. The average entropy production is then the expectation value of this operator over the physical process. Notably, this definition automatically satisfies the Jarzynski equality.
The two definitions are not identical in general, and we illustrate this using a canonical family of channels (Section V). In the conclusion (Section VI), we summarize where these results leave us in the quest to find a systematic definition of entropy production, valid for all physical processes.
II Classical entropy production and Bayesian retrodiction
In classical stochastic thermodynamics [7, 8, 9, 10], physical processes are described by the joint probability distributions of the input states $s_{i}$ and the output states $s_{o}$. The ingredients are the distribution $p(s_{i})$ of initial states, that can be chosen, at least in principle; and the physical (stochastic) evolution $\varphi(s_{o}|s_{i})$, dictated by the laws of physics and our limitations (notably coarse-graining of the microstates into macrostates). The probability of each physical trajectory $s_{i}\to s_{o}$ is then computed as
$P_{F}(s_{i},s_{o})$ $\equiv\,p(s_{i})\varphi(s_{o}|s_{i})\;.$ (1)
The above is also known as a forward process, and it is countered by a corresponding reverse process. The construction of the latter can be considered intuitive in some cases, and has been derived from physical considerations in others; but only recently has a general recipe been proposed [11, 12] that recovers the previous ones as special cases. This recipe treats the reverse process as a Bayesian retrodiction performed with respect to a prior distribution $\pi(s_{i})$:
$\hat{\varphi}_{\pi}(s_{i}|s_{o})$ $\equiv\varphi(s_{o}|s_{i})\frac{\pi(s_{i})}{(\varphi\pi)(s_{o})}\,,$ (2)
where the notation $(\varphi p)$ refers to the posterior of $p$ after the channel $\varphi$ has been applied, i.e., $(\varphi p)(s_{o})\equiv\sum_{s_{i}}p(s_{i})\varphi(s_{o}|s_{i})$. The choice of the prior $\pi$ on which the Bayesian inversion is performed typically depends on the situation: for an adiabatic process, it is often chosen to be the uniform prior, while for an isothermal process, it is often chosen to be the Gibbs distribution at the given temperature. More generally, a common choice is to have $\pi$ as the fixed point for the forward channel. Note, however, that the choice of $\pi$ can be quite arbitrary: in particular, it need not coincide with the “true” initial distribution $p(s_{i})$. Another point to note is that the Bayesian definition of the reverse process is equivalent to the one obtained by “dilation”, i.e. the recipe: add a reservoir to close the system, invert the deterministic trajectories of the closed system, then remove the reservoir [13].
Having defined the reverse evolution as in (2), one can choose its initial distribution $q(s_{o})$. As it was the case for $\pi$, this also can be chosen quite freely: in particular, it may or may not be the same as the final distributions $(\varphi p)$ or $(\varphi\pi)$ of the forward process. Then, the probability of the reverse trajectory $s_{o}\to s_{i}$ is
$P_{R}^{\pi}(s_{o},s_{i})$ $=q(s_{o})\,\hat{\varphi}_{\pi}(s_{i}|s_{o})\;.$ (3)
The irreversible entropy production on the stochastic trajectory $s_{i}\rightarrow s_{o}$ then becomes
$\sigma[P_{F},P_{R}^{\pi}](s_{i},s_{o})$ $\equiv\log\frac{P_{F}(s_{i},s_{o})}{P_{R}(s_{o},s_{i})}$ (4)
$=\log\frac{p(s_{i})}{\pi(s_{i})}-\log\frac{(\varphi\pi)(s_{o})}{q(s_{o})}\,.$ (5)
The first line is the usual definition; from it, the Crooks fluctuation theorem [14] and Jarzynski equality [15] follow as formal identities [16, 11, 12]. The second line reveals another feature proper to the Bayesian inversion recipe (see Result 3 in [12]): the entropy production becomes the sum of a function of $s_{i}$ and a function of $s_{o}$ if and only if the reverse process follows the Bayesian recipe. Note that this does not mean that entropy production is a state function: for that to be the case, the two functions would have to be the same, which is generally not the case here. Nevertheless, this fact, together with the dilation argument recalled earlier, gives the Bayesian approach a special place among all possible choices of reversal.
The average entropy production of the physical process becomes
$\left\langle\sigma\right\rangle_{F}$ $=\,D(P_{F}\|P_{R}^{\pi})$ (6)
$=D_{\textrm{in}}(p\|\pi)-D_{\textrm{out}}(\varphi p\|\varphi\pi)+D_{\textrm{out}}(\varphi p\|q)$ (7)
where $D$ is the Kullback–Leibler (KL) divergence [17], and where the subscripts in the second line stress how the expression separates between input and output. Again, the first line is a well-known expression, first noticed in Ref. [18], while its expansion in the second line (7) relies on expressing the reverse process as a Bayesian retrodiction. To the best of our knowledge, this is the first time it has been written down explicitly at this level of generality, although special cases have often been used. For example, the version without the last term (i.e. for $q=\varphi p$) was derived from more fundamental considerations, e.g. in Ref. [19]; and the further case where $\pi$ is the Gibbs state, usually assumed to satisfy $\pi=\varphi\pi$, is quite common in thermodynamics, useful for directly linking entropy production to the change in non-equilibrium free energy [20].
Our goal is to extend these considerations to fully quantum expressions, with probability distributions replaced by (generally non-commuting) quantum states.
III First definition
As it turns out, the average $\left\langle\sigma\right\rangle_{F}$ as expressed in Eq. 7 can be immediately extended to quantum theory, as the first term depends only on the input states, and the second and third only on the output states. Thus, for arbitrary initial state $\rho$, reference state $\gamma$, input of the reverse process $\tau$, and CPTP map $\mathcal{E}$, one could declare its fully quantum version to be
$\left\langle\Sigma\right\rangle$ =$\stackrel{{\scriptstyle\textrm{def.~{}1}}}{{=}}\,D(\rho\|\gamma)-D(\mathcal{E}(\rho)\|\mathcal{E}(\gamma))+D(\mathcal{E}(\rho)\|\tau)$ (8)
where $D$ is Umegaki’s relative entropy [21, 22]. This expression fulfills at least two basic desiderata: it is always non-negative, since $D(\rho\|\gamma)\geqslant D(\mathcal{E}(\rho)\|\mathcal{E}(\gamma))$ and $D(\mathcal{E}(\rho)\|\tau)\geqslant 0$; and it coincides with the classical expression (7) when everything commutes.
If we were to stop our analysis here, we would already have a definition of irreversible entropy production for quantum processes that recovers and generalizes several previous formulations [23, 2], different from the eleven physical versions studied by Strasberg and Schindler [3] and from the three proposals presented in Ref. [6].
However, this approach would forgo two key features of the classical formulation:
  • We would only have an analog of the average $\left\langle\sigma\right\rangle_{F}$, but without knowing what is being averaged, or the specific distribution over which it is averaged—i.e., without knowing whether there is a fully quantum analog of the detailed quantity $\sigma$. Knowledge of $\sigma$ is essential for calculating further statistical properties, notably the Crooks fluctuation theorem [14] and the Jarzynski equality $\left\langle e^{-\sigma}\right\rangle_{F}=1$ [15].
  • Although in (8) the state $\gamma$ would still play the role of “prior belief”, the connection with Bayesian inference would be significantly weakened—if not entirely lost—as there would be no direct reference to a statistical comparison between a forward process and its reverse as in (6).
So, in a positive way, if we could find a stochastic entropy production operator $\Sigma$, we could derive fully quantum fluctuation theorems by studying its algebraic and spectral properties directly, instead of relying on semi-classical ad hoc solutions like the two-measurement approach and its generalizations [24, 25, 4]. And if this operator involves a “quantum forward process” and its associated “quantum reverse process”, we would retain the strong connection with logical inference provided by retrodiction. The mathematical challenge is that, while joint input-output probability distributions are commonplace in classical theory, such two-time objects are much less obvious in quantum theory.
In what follows, we fill this gap by defining a two-time quantum object $Q_{F}$ describing the input-output process, and constructing its reverse $Q_{R}^{\gamma}$ using a Bayesian recipe. We then construct a non-commuting counterpart $\Sigma[Q_{F},Q_{R}^{\gamma}]$ of (4), which is a stochastic entropy production operator. We then study the properties of this operator and its average. To the best of our knowledge, this is the first time such an approach has been attempted.
IV Second definition: an entropy production operator
There is a growing interest in defining objects in the quantum formalism, called states-over-time, that encode both the input $\rho$ and the output $\mathcal{E}(\rho)$ of a process (described by a CPTP linear map $\mathcal{E}$) [26, 27, 28, 29, 30]. Different definitions satisfy different desiderata. In our context, we are looking for a generalization of joint probability distributions: the only desideratum imposed by this analogy is that the object $Q$, which we use as the quantum analog of the classical trajectory probability distribution, be a density operator, i.e., a non-negative Hermitian operator normalized so that $\mathrm{Tr}[Q]=1$. We would thus have one such object for the forward process, i.e., $Q_{F}$, and another for the reverse, i.e., $Q^{\gamma}_{R}$, where $\gamma$ is the quantum state (analogous to the prior $\pi$) with respect to which the reverse is constructed.
Other desiderata concern the properties of the microscopic entropy production operator $\Sigma[Q_{F},Q_{R}^{\gamma}]$ constructed from such objects. At the very least, we want its average to be non-negative, so that the Second Law retains its usual formulation
$\left\langle\Sigma[Q_{F},Q_{R}^{\gamma}]\right\rangle_{F}\equiv\mathrm{Tr}\Big{\{}Q_{F}\;\Sigma[Q_{F},Q_{R}^{\gamma}]\Big{\}}\geqslant 0\;.$
Note that some of the quantum entropy production proposed in [6] do not satisfy this condition, as the average may take complex values. Almost as strongly, we would like $\Sigma[Q_{F},Q_{R}^{\gamma}]$ to be Hermitian, so that the individual values that the entropy can take are also real. An example where the entropy can take complex values, even if the average turns out to be real and non-negative, can be found in [4].
Besides that, we think that the “local” character of Bayesian entropy production manifested in Eq. 5 is a deep feature, and we hope to retain it in some form. Finally, its average $\left\langle\Sigma\right\rangle_{F}$ should of course recover the classical expression when describing a classical process, and possibly preserve the structure (7) in other situations as well.
IV.1 Definition
We start by defining a two-time quantum object corresponding to the forward process, as follows
$\tilde{Q}_{F}(\rho)$ $\equiv(\mathds{1}\otimes\sqrt{\rho^{T}})\,C_{\mathcal{E}}\,(\mathds{1}\otimes\sqrt{\rho^{T}}).$ (9)
where $C_{\mathcal{E}}\equiv\sum_{i,j}\mathcal{E}(|i\rangle\!\!\!\;\langle j|)\otimes|i\rangle\!\!\!\;\langle j|$ is the Choi operator [31] for channel $\mathcal{E}$, defined with a basis {|$\{\ket{i}\}$ of the input system. This operator $\tilde{Q}_{F}(\rho)$ is non-negative and unit-trace: it is thus a valid bipartite state. Moreover, it encodes the initial and final states, which can be recovered as $\textrm{Tr}_{A}[\tilde{Q}_{F}(\rho)]=\rho^{T}$ and $\textrm{Tr}_{B}[\tilde{Q}_{F}(\rho)]=\mathcal{E}(\rho)$. Such a two-time object has appeared in studies of quantum conditional probability [32, 33] and observational entropy [34], and is similar to Leifer–Spekkens state over time [26, 27].
Next, we introduce the corresponding reverse process. Following the Bayesian prescription discussed above, we define the reverse process with respect to a prior state $\gamma$ using the Petz transpose map [35, 36]:
$\mathcal{R}_{\mathcal{E}}^{\gamma}(\tau)\equiv\sqrt{\gamma}\,\mathcal{E}^{\dagger}\left[\mathcal{E}(\gamma)^{-1/2}\tau\mathcal{E}(\gamma)^{-1/2}\right]\sqrt{\gamma}\,.$ (10)
In the above equation, $\tau$ is the input for the reverse channel modeling the new (quantum) information, analogous to the distribution $q(s_{o})$ in Eq. 3. The choice of the Petz transpose map as the quantum Bayes rule in the work is justified in light of what we discuss in Ref. [37]. There we show how the Petz transpose map emerges as the least committal update in the face of new information, that is, the update rule that changes the prior belief as little as possible, compatibly with the new information. This is strictly true only when $[\tau,\mathcal{E}(\gamma)]=0$, but this is also the situation in which the entropy production operator defined here enjoys the best mathematical properties  [^1] Otherwise, if $[\tau,\mathcal{E}(\gamma)]\neq 0$, the optimal update rule that we obtain in [37] is a variant of Petz’s map.. For the sake of the present discussion, therefore, we will only work with the Petz transpose map, as we already did in other works [11, 38, 39].
As shown in [34], the Choi operator of $\mathcal{R}_{\mathcal{E}}^{\gamma}$ can be related to that of $\mathcal{E}$ as follows:
$C_{\mathcal{R}_{\mathcal{E}}^{\gamma}}^{T}$
$=(\sqrt{\tau}\mathcal{E}(\gamma)^{-1/2}\otimes\sqrt{\gamma^{T}})\,C_{\mathcal{E}}\,(\mathcal{E}(\gamma)^{-1/2}\sqrt{\tau}\otimes\sqrt{\gamma^{T}})\;.$
The corresponding two-time object is then defined similarly to Eq. 9 as follows:
$\tilde{Q}_{R}^{\gamma}(\tau)$
$\equiv(\sqrt{\tau}\otimes\mathds{1})C_{\mathcal{R}_{\mathcal{E}}^{\gamma}}^{T}(\sqrt{\tau}\otimes\mathds{1})$
$=(\sqrt{\tau}\mathcal{E}(\gamma)^{-1/2}\otimes\sqrt{\gamma^{T}})\,C_{\mathcal{E}}\,(\mathcal{E}(\gamma)^{-1/2}\sqrt{\tau}\otimes\sqrt{\gamma^{T}})\,.$
In fact, as the analogs of $P_{F}$ and $P_{R}^{\pi}$ we will not directly use $\tilde{Q}_{F}$ and $\tilde{Q}_{R}^{\gamma}$, but rather their pseudotransposes defined as
$Q_{F}(\rho)$ $\equiv\sqrt{C_{\mathcal{E}}}\,(\mathds{1}\otimes\rho^{T})\,\sqrt{C_{\mathcal{E}}}\,,$ (11)
$Q_{R}^{\gamma}(\tau)$ $\equiv\sqrt{C_{\mathcal{E}}}\left(\mathcal{E}(\gamma)^{-1/2}\tau\mathcal{E}(\gamma)^{-1/2}\otimes\gamma^{T}\right)\sqrt{C_{\mathcal{E}}}\,.$ (12)
These operators, similarly to $\tilde{Q}_{F}$ and $\tilde{Q}_{R}^{\gamma}$, are valid bipartite states, but in addition their dependence on $\rho$ and $\tau$ is linear. In fact, as shown in Appendix A, $Q_{F}$ and $Q_{R}^{\gamma}$ are both well-defined quantum channels acting on $\rho$ and on $\tau$, respectively. More precisely, $Q_{F}$ is equivalent to the complementary channel of $\mathcal{E}$, up to a transpose. Similarly, $Q_{R}^{\gamma}$ is equivalent to the complementary channel of the reverse map $\mathcal{R}_{\mathcal{E}}^{\gamma}$.
We are now in the position to introduce, after Eq. 8, our second definition of entropy production, this time not just as an average but as the stochastic entropy-production operator:
$\Sigma[Q_{F}(\rho),Q_{R}^{\gamma}(\tau)]$ (13)
=$\stackrel{{\scriptstyle\textrm{def.~{}2}}}{{=}}\log\left\{\sqrt{Q_{F}(\rho)}\Big{[}Q_{R}^{\gamma}(\tau)\Big{]}^{-1}\sqrt{Q_{F}(\rho)}\right\}\,.$
From the above definition, it is clear that we must restrict ourselves to those situations where $Q_{R}^{\gamma}(\tau)$ is invertible on the support of $Q_{F}(\rho)$, so as not to run into division by zero. This is a ubiquitous “problem” when dealing with inversions. The logarithm must also be restricted to the support of its argument. Nevertheless, when considering the average entropy production, it is possible to treat rank-deficient cases (most importantly the case of unitary channels) on a case-by-case basis, as we will do below in Eq. 25. Finally, for the sake of readability, we will refrain from explicitly writing the dependence on $\rho$ and $\tau$ from here on, where there is no confusion, and denote $Q_{F}(\rho)$ and $Q_{R}^{\gamma}(\tau)$ as $Q_{F}$ and $Q_{R}^{\gamma}$, respectively.
IV.2 Basic properties of the operator
As argued above, the main desideratum for us is that the average entropy production should be non-negative. This is satisfied by construction, since
$\left\langle\Sigma[Q_{F},Q_{R}^{\gamma}]\right\rangle_{F}$ $\equiv\mathrm{Tr}\Big{\{}Q_{F}\;\Sigma[Q_{F},Q_{R}^{\gamma}]\Big{\}}$
$=\textrm{Tr}\left[Q_{F}\log\sqrt{Q_{F}}(Q_{R}^{\gamma})^{-1}\sqrt{Q_{F}}\right]$
$\equiv D_{\textrm{BS}}(Q_{F}\|Q_{R}^{\gamma})\,\geqslant 0$ (14)
where $D_{\textrm{BS}}$ is the Belavkin–Staszewski relative entropy [40, 41]. This relative entropy is an upper bound on Umegaki’s relative entropy and, for commuting operators, coincides with the KL-divergence. We refer the reader to Refs. [42, 43] for its further properties and relations to other divergences. Indeed, any valid relative entropy between $Q_{F}$ and $Q_{R}^{\gamma}$ is going to be non-negative, since they are non-negative unit-trace operators. Moreover, $D_{\textrm{BS}}(Q_{F}\|Q_{R}^{\gamma})=0$ if and only if $Q_{F}=Q_{R}^{\gamma}$, which, in the case of full rank $C_{\mathcal{E}}$, occurs if and only if $\rho=\gamma$ and $\tau=\mathcal{E}(\rho)=\mathcal{E}(\gamma)$. In this case, our prior belief $\gamma$ was “correct” and no new information was acquired. We shall expand on the explicit expression of (14) later.
IV.3 Fully quantum fluctuation theorems
Just as classical fluctuation theorems are formal consequences of the definition Eq. 4 of entropy production, so the formal properties of the operator (13) can be considered as fully quantum fluctuation theorem. It is easy to verify that that operator always satisfies the fully quantum Jarzynski equality:
$\left\langle e^{-\Sigma[Q_{F},Q_{R}^{\gamma}]}\right\rangle_{F}$ $=\textrm{Tr}\left[Q_{F}\;e^{-\Sigma[Q_{F},Q_{R}^{\gamma}]}\right]$
$=\mathrm{Tr}[Q_{R}^{\gamma}]=1\,,$ (15)
whenever $Q_{F}$ and $Q_{R}^{\gamma}$ are supported in the same subspace, with the same formal meaning as in the classical case, namely the normalization of the reverse process.
By contrast, the generalisation of Crooks’ fluctuation theorem shows a difference between classical and quantum fluctuations in the case of non-commuting processes. Since our operator is Hermitian, we can define the values of entropy production as its eigenvalues:
$\Sigma[Q_{F},Q_{R}^{\gamma}]\ket{f_{k}}$ $=\,\Sigma_{k}\ket{f_{k}}\,.$ (16)
Assuming all the eigenvalues are non-degenerate, the distribution of these values in the forward process is given by
$\mathbb{P}_{F}(\Sigma_{k})$ $=\,\langle f_{k}|Q_{F}|f_{k}\rangle\,.$ (17)
Next we need to define entropy production in the reverse process. The operator is obviously given by exchanging the roles of $Q_{R}^{\gamma}$ and $Q_{F}$. Contrary to the classical case, the identity $\Sigma[Q_{R}^{\gamma},Q_{F}]=-\Sigma[Q_{F},Q_{R}^{\gamma}]$ does not hold in general. The eigenvalues are indeed the same with opposite signs, as can be proved notcing that $e^{\Sigma[Q_{F},Q_{R}^{\gamma}]}=\sqrt{Q_{F}}(Q_{R}^{\gamma})^{-1}\sqrt{Q_{F}}$ and $e^{-\Sigma[Q_{R}^{\gamma},Q_{F}]}=(Q_{R}^{\gamma})^{-1/2}Q_{F}(Q_{R}^{\gamma})^{-1/2}$ have the same eigenvalues by polar decomposition. The eigenvectors are in general different, and so
$\Sigma[Q_{R}^{\gamma},Q_{F}]\ket{r_{k}}$ $=\,-\Sigma_{k}\ket{r_{k}}\,.$ (18)
The statistics of entropy on the reverse process are thus given by
$\mathbb{P}_{R}(-\Sigma_{k})$ $=\,\langle r_{k}|Q_{R}^{\gamma}|r_{k}\rangle=\,\sum_{j}e^{-\Sigma_{j}}|\langle r_{k}|\sqrt{Q_{F}}|f_{j}\rangle|^{2}\,,$ (19)
where we have rewritten Eq. 13 as $Q_{R}^{\gamma}=\sqrt{Q_{F}}e^{-\Sigma[Q_{F},Q_{R}^{\gamma}]}\sqrt{Q_{F}}$. Eq. 19 is the fully quantum extension of the Crooks theorem. The classical Crooks expression
$\mathbb{P}_{R}(-\Sigma_{k})$ $=\,e^{-\Sigma_{k}}\,\mathbb{P}_{F}(\Sigma_{k})$ (20)
is recovered only when $[Q_{F},Q_{R}^{\gamma}]=0$. In the general case, non-commutativity (i.e. coherence) impacts the fluctuations of entropy. Still, like in the classical case, Jarzynski equality (IV.3) can be derived starting from Eq. 19, as $\sum_{k}\mathbb{P}_{R}(-\Sigma_{k})=\sum_{j}e^{-\Sigma_{j}}\mathbb{P}_{F}(\Sigma_{j})=1$. A comparison between the quantum and classical expressions is shown in Fig. 1 for the channel that we shall use in Section V. One might surmise that quantum fluctuations are always larger than classical ones. The figure shows that this is often the case, but not always: for some $\rho$, the standard deviation of the quantum distribution is actually smaller.
Figure 1: Comparison between the statistics of entropy in the quantum Eq. 19 and the classical Crooks expression Eq. 20 for a one-qubit channel (the entropy production operator is of dimension 4). The channel $\mathcal{N}$ is defined by Eq. 35 where $\gamma$ is chosen to be equal to $\xi$, and $\tau=\mathcal{N}(\rho)$. For the numerics, we choose $\langle 0|\xi|0\rangle=0.8$, $\phi=0.8$ and $n=1$. The plot represents the $(x,z)$ plane of the Bloch sphere of $\rho$. The heat map represents the standard deviation of the reverse distribution $\mathbb{P}_{R}$: left, for Eq. 19 obtained from studying our operator; right, computed from $\mathbb{P}_{F}$ assuming the classical Crooks relation Eq. 20. For most $\rho$, the quantum expression shows a larger standard deviation than the classical guess; but for some, the opposite holds.
IV.4 Locality in time and the Heisenberg work operator
As shown in Eq. 5, the classical entropy production $\sigma[P_{F},P_{R}^{\pi}]$ can be decomposed into two terms, one depending only on the initial state, the other only on the final state. We call this property locality in time. For the entropy production operator $\Sigma[Q_{F},Q_{R}^{\gamma}]$, when the Choi operator $C_{\mathcal{E}}$ is full rank, we have
$U\;\Sigma[Q_{F},Q_{R}^{\gamma}]\;U^{\dagger}$ $=\mathds{1}\otimes\log\sqrt{\rho^{T}}(\gamma^{T})^{-1}\sqrt{\rho^{T}}$
$-\log\mathcal{E}(\gamma)^{-1/2}\tau\mathcal{E}(\gamma)^{-1/2}\otimes\mathds{1}\;.$ (21)
where $U\equiv(\mathds{1}\otimes\rho^{T})^{-1/2}C_{\mathcal{E}}^{-1/2}\sqrt{Q_{F}}$ is a global unitary  [^2] In case $\rho$ is rank-deficient, one can define $U$ with the polar decomposition of $(\mathds{1}\otimes\rho^{T})C_{\mathcal{E}}$ as $(\mathds{1}\otimes\rho^{T})C_{\mathcal{E}}=U\sqrt{Q_{F}}$.. Thus locality in time does not hold generically for the operator, although it does for its eigenvalues. If $[(\mathds{1}\otimes\rho^{T}),C_{\mathcal{E}}]=0$ is satisfied (notably, for classical-quantum channels), then $U=\mathds{1}$ and locality in time holds exactly.
In the special case, common in the context of quantum thermodynamics, in which the prior state $\gamma$ is uniform, $\mathcal{E}$ is unital, the initial state of the system $\rho$ is thermal with respect to the initial Hamiltonian $H$, and the initial state of the reverse process $\tau$ is thermal with respect to the final Hamiltonian $H^{\prime}$, Section IV.4 reduces to the familiar expression
$U\;\Sigma[Q_{F},Q_{R}^{\gamma}]\;U^{\dagger}$ $=\beta(\mathds{1}\otimes H^{\prime}-H\otimes\mathds{1}-\Delta F)$
$=\beta(\Omega-\Delta F)\;,$
where $\Omega$ is the so-called Heisenberg operator of work [44, 45]. We leave the exploration of further connections between our entropy production operator and the Heisenberg work operator to future investigations.
Moving on to the average entropy production Eq. 14, the relation simply further once we notice that $UQ_{F}U^{\dagger}=\tilde{Q}_{F}$, with $\tilde{Q}_{F}$ defined in Eq. 9. Therefore, for all well-behaved (analytical) function $f$ such as, e.g., the moment generating function, we have
$\textrm{Tr}\left[Q_{F}\ f(\Sigma[Q_{F},Q_{R}^{\gamma}])\right]$
$=\mathrm{Tr}\left[Q_{F}\ U^{\dagger}U\ f(\Sigma[Q_{F},Q_{R}^{\gamma}])\ U^{\dagger}U\right]$
$=\mathrm{Tr}\left[UQ_{F}U^{\dagger}\ f(U\Sigma[Q_{F},Q_{R}^{\gamma}]U^{\dagger})\right]$
$=\mathrm{Tr}\Big{[}\tilde{Q}_{F}\ f\big{(}\mathds{1}\otimes\log\sqrt{\rho^{T}}(\gamma^{T})^{-1}\sqrt{\rho^{T}}$
$\mathrel{\phantom{=}}\qquad-\log\mathcal{E}(\gamma)^{-1/2}\tau\mathcal{E}(\gamma)^{-1/2}\otimes\mathds{1}\big{)}\Big{]}\;.$ (22)
IV.5 Explicit expressions for the average entropy production
We turn now to study more closely the average entropy production (14). For channels with full-rank Choi operator, we find the explicit expression
$\left\langle\Sigma[Q_{F},Q_{R}^{\gamma}]\right\rangle_{F}=$
$D_{\textrm{BS}}(\rho\|\gamma)-\textrm{Tr}\left[\mathcal{E}(\rho)\log\left(\mathcal{E}(\gamma)^{-1/2}\tau\mathcal{E}(\gamma)^{-1/2}\right)\right]\,,$ (23)
which can be derived using Eq. 22. When $[\mathcal{E}(\gamma),\tau]=0$, this reduces to (8), up to replacing the Umegaki relative entropy with the Belavkin–Staszewski one for the input term:
$\left\langle\Sigma[Q_{F},Q_{R}^{\gamma}]\right\rangle_{F}\stackrel{{\scriptstyle[\mathcal{E}(\gamma),\tau]=0}}{{=}}$
$D_{\textrm{BS}}(\rho\|\gamma)-D(\mathcal{E}(\rho)\|\mathcal{E}(\gamma))+D(\mathcal{E}(\rho)\|\tau)\,.$ (24)
Note that the commutation condition required for the above to hold involves only the prior $\gamma$ and the initial state of the reverse process $\tau$, both of which can be chosen independently of the “true” initial state of the system $\rho$. In particular, the “true” output $\mathcal{E}(\rho)$ need not commute with the new quantum information $\tau$. A special case in which $[\mathcal{E}(\gamma),\tau]=0$ is always satisfied is that of channels with commutative output, as in the case of measurements. See Section IV.7 below.
We do not investigate every possible rank-deficient channel, but only the important case of unitary channels $\mathcal{U}$. In this case, the Choi operator being a rank-one projector, $Q_{R}^{\gamma}(\tau)=Q_{F}(\rho)$ holds for all choices of states $\rho$, $\gamma$, and $\tau$. Therefore,
$\left\langle\Sigma[Q_{F},Q_{R}^{\gamma}]\right\rangle_{F}$ =$\stackrel{{\scriptstyle\mathcal{E}=\mathcal{U}}}{{=}}0\,.$ (25)
The fact that no entropy is produced by a unitary channel is certainly expected.
Finally, we notice that in many studies the start of the reverse process is taken as the output of the forward process, i.e. $\tau=\mathcal{E}(\rho)$. By imposing only this condition, we do not find any simplifications of Section IV.5, but we have the inequality
$\left\langle\Sigma[Q_{F},Q_{R}^{\gamma}]\right\rangle_{F}\stackrel{{\scriptstyle\tau=\mathcal{E}(\rho)}}{{\geqslant}}D_{\textrm{BS}}(\rho\|\gamma)$ $-D(\mathcal{E}(\rho)\|\mathcal{E}(\gamma))\,.$ (26)
The proof is in Appendix B. The equality does not hold in general, but it does hold if we further have $[\mathcal{E}(\gamma),\tau]\equiv[\mathcal{E}(\gamma),\mathcal{E}(\rho)]=0$, because then it follows directly from (24).
IV.6 Superadditivity and mutual entropy production
The average entropy production satisfies a superadditivity property. Let us consider two channels, $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$, that can be composed in series, so that the composition $\mathcal{E}_{2}\circ\mathcal{E}_{1}=:\mathcal{E}$ is well-defined. Let the true initial state be $\rho$, $\gamma$ be the prior used to construct the reverse of $\mathcal{E}_{1}$, and $\mathcal{E}_{1}(\gamma)$ be the prior used to construct the reverse of $\mathcal{E}_{2}$. Finally, let $\tau_{1}$ be the new information obtained at the output of channel $\mathcal{E}_{1}$, and $\tau_{2}$ the new information obtained at the output of channel $\mathcal{E}_{2}$.
The two-time operators (forward and reverse) for the whole process $\mathcal{E}=\mathcal{E}_{2}\circ\mathcal{E}_{1}$ are given by
$Q_{F}(\rho)\equiv\sqrt{C_{\mathcal{E}}}(\mathds{1}\otimes\rho^{T})\sqrt{C_{\mathcal{E}}}$
$Q_{R}^{\gamma}(\tau_{2})\equiv$
$\quad\sqrt{C_{\mathcal{E}}}\left(\mathcal{E}(\gamma)^{-1/2}\tau_{2}\mathcal{E}(\gamma)^{-1/2}\otimes\mathcal{E}_{1}(\gamma)^{T}\right)\!\sqrt{C_{\mathcal{E}}}\;.$
The corresponding operators for the two steps separately are given by
$Q_{F,1}(\rho)\equiv\sqrt{C_{\mathcal{E}_{1}}}(\mathds{1}\otimes\rho^{T})\sqrt{C_{\mathcal{E}_{1}}}$
$Q_{R,1}^{\gamma}(\tau_{1})\equiv$
$\quad\sqrt{C_{\mathcal{E}_{1}}}\left(\mathcal{E}_{1}(\gamma)^{-1/2}\tau_{1}\mathcal{E}_{1}(\gamma)^{-1/2}\otimes\gamma^{T}\right)\sqrt{C_{\mathcal{E}_{1}}}\;,$
and
$Q_{F,2}(\mathcal{E}_{1}(\rho))\equiv\sqrt{C_{\mathcal{E}_{2}}}\left(\mathds{1}\otimes\mathcal{E}_{1}(\rho)^{T}\right)\sqrt{C_{\mathcal{E}_{2}}}$
$Q_{R,2}^{\mathcal{E}_{1}(\gamma)}(\tau_{2})\equiv$
$\quad\sqrt{C_{\mathcal{E}_{2}}}\left(\mathcal{E}(\gamma)^{-1/2}\tau_{2}\mathcal{E}(\gamma)^{-1/2}\otimes\mathcal{E}_{1}(\gamma)^{T}\right)\sqrt{C_{\mathcal{E}_{2}}}\,.$
Notice that the reverse process for the second step is defined with respect to the reference prior $\mathcal{E}_{1}(\gamma)$.
The entropy production operators for the first step, the second step, and the total process are then defined as follows:
$\Sigma_{1}$ $\equiv\Sigma[Q_{F,1}(\rho),Q_{R,1}^{\gamma}(\tau_{1})]$
$\Sigma_{2}$ $\equiv\Sigma[Q_{F,2}(\mathcal{E}_{1}(\rho)),Q_{R,2}^{\mathcal{E}_{1}(\gamma)}(\tau_{2})]$
$\Sigma_{12}$ $\equiv\Sigma[Q_{F}(\rho),Q_{R}^{\gamma}(\tau_{2})]\,,$
their averages being
$\left\langle\Sigma_{1}\right\rangle_{F,1}$ $\equiv\mathrm{Tr}[Q_{F,1}(\rho)\ \Sigma_{1}]$
$\left\langle\Sigma_{2}\right\rangle_{F,1}$ $\equiv\mathrm{Tr}[Q_{F,2}(\mathcal{E}_{1}(\rho))\ \Sigma_{2}]$
$\left\langle\Sigma_{12}\right\rangle_{F}$ $\equiv\mathrm{Tr}[Q_{F}(\rho)\ \Sigma_{12}]\;.$
With the above definitions, for full-rank $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$, we have
$\left\langle\Sigma_{1}\right\rangle_{F,1}+\left\langle\Sigma_{2}\right\rangle_{F,2}\geqslant\left\langle\Sigma_{12}\right\rangle_{F}\,,$ (27)
for any choice of $\tau_{1}$ and $\tau_{2}$. Equality holds if and only if $\tau_{1}=\mathcal{E}_{1}(\rho)$ and $[\mathcal{E}_{1}(\rho),\mathcal{E}_{1}(\gamma)]=0$. The proof of this property is in Appendix C.
It is tempting to interpret the above inequality as a time-like analog of the positivity of mutual information, as Eq. 27 could be rewritten as
$\left\langle\Upsilon_{1:2}\right\rangle_{F}\equiv\left\langle\Sigma_{1}\right\rangle_{F,1}+\left\langle\Sigma_{2}\right\rangle_{F,2}-\left\langle\Sigma_{12}\right\rangle_{F}\geqslant 0\;.$
We leave it for future investigation whether the above is merely a formal analogy or something more.
IV.7 Quantum-classical channels
Quantum-classical channels are channels with fully quantum input but commutative output. This type of channels can be understood as describing measurements done on the input system, with the outcome probabilities encoded in the diagonal elements of the output:
$\mathcal{M}(\rho)=\sum_{i}\mathrm{Tr}[\Pi_{i}\rho]|i\rangle\!\!\!\;\langle i|\;,$ (28)
for a POVM $\{\Pi_{i}\}_{i}$ and an orthonormal basis {|$\{\ket{i}\}_{i}$. In this setup, by construction, $\mathcal{M}(\rho),\mathcal{M}(\gamma)$ are diagonal in the basis {|$\{\ket{i}\}$, and it is natural to assume that also $\tau$ be diagonal in the same basis.
As a consequence, in the quantum-classical case, the operators $C_{\mathcal{M}},\mathcal{M}(\rho)\otimes\mathds{1},\mathcal{M}(\gamma)\otimes\mathds{1},\tau\otimes\mathds{1}$ all commute, and the entropy production operator has a block-diagonal form:
$\Sigma[Q_{F},Q_{R}^{\gamma}]=$ $\sum_{i}|i\rangle\!\!\!\;\langle i|\otimes\Sigma\left[\sqrt{\Pi_{i}}\rho\sqrt{\Pi_{i}},\sqrt{\Pi_{i}}\gamma\sqrt{\Pi_{i}}\right]^{T}$
$\mathrel{\phantom{=}}-(\ln\tau-\ln\mathcal{M}(\gamma))\otimes\mathds{1}\;,$ (29)
where
$\Sigma\left[\sqrt{\Pi_{i}}\rho\sqrt{\Pi_{i}},\sqrt{\Pi_{i}}\gamma\sqrt{\Pi_{i}}\right]$
$=\log\sqrt{\sqrt{\Pi_{i}}\rho\sqrt{\Pi_{i}}}\Big{(}\sqrt{\Pi_{i}}\gamma\sqrt{\Pi_{i}}\Big{)}^{-1}\sqrt{\sqrt{\Pi_{i}}\rho\sqrt{\Pi_{i}}}\;.$ (30)
The average entropy production becomes
$\left\langle\Sigma[Q_{F},Q_{R}^{\gamma}]\right\rangle_{F}=D_{\textrm{BS}}(\rho\|\gamma)$ $-D(\mathcal{M}(\rho)\|\mathcal{M}(\gamma))$
$+D(\mathcal{M}(\rho)\|\tau)\,,$ (31)
where the last two terms are in fact relative entropies of classical distributions, since all those states are diagonal in the same basis.
This construction tallies exactly with a recent definition of observational entropy with general quantum priors [34]. Observational entropy $S_{\mathsf{M}}(\rho)$ is defined in the context of an observation, i.e. a measurement $\mathsf{M}=\{\Pi_{i}\}_{i}$: it captures total entropy as the sum of the intrinsic uncertainty of the system $S(\rho)$ and of the additional uncertainty induced by the measurement, which may not be maximally informative [38]. For quantum prior knowledge, it was proposed that observational entropy generalizes as [34]
$S_{\mathsf{M}}(\rho)=S(\rho)+D_{\textrm{BS}}(\rho\|\gamma)$ $-D(\mathcal{M}(\rho)\|\mathcal{M}(\gamma))\,.$ (32)
Since in this context, the new information about the outcome is given by the statistics of the measurement, i.e., $\tau=\mathcal{M}(\rho)$, we see from Eq. 31 that the entropy $S_{\mathsf{M}}(\rho)-S(\rho)$ induced by the observation is exactly the entropy production $\left\langle\Sigma[Q_{F},Q_{R}^{\gamma}]\right\rangle_{F}$ of the measurement process.
IV.8 Classical case
The classical case is recovered when all the states are diagonal in one basis, and the channel is the measure-and-prepare channel on the same basis. In this case, the operators $C_{\mathcal{E}},\mathds{1}\otimes\rho^{T},\mathds{1}\otimes\gamma^{T},\mathcal{E}(\rho)\otimes\mathds{1},\mathcal{E}(\gamma)\otimes\mathds{1},\tau\otimes\mathds{1}$ all commute. The entropy production operator thus simplifies to
$\Sigma[Q_{F},Q_{R}^{\gamma}]$
$=\mathds{1}\otimes(\log\rho-\log\gamma)-\mathopen{\big{(}}\log\tau-\log\mathcal{E}(\gamma)\big{)}\otimes\mathds{1}$ (33)
$=\sum_{i,j}\log\frac{p_{i}\,\lambda_{j}}{\pi_{i}\,q_{j}}\,|j\rangle\!\!\!\;\langle j|\otimes|i\rangle\!\!\!\;\langle i|$ (34)
where $p$, $\pi$, $\lambda=\varphi\pi$ and $q$ are the eigenvalues of $\rho$, $\gamma$, $\mathcal{E}(\gamma)$ and $\tau$, and $\varphi$ is the stochastic matrix induced by $\mathcal{E}$ on the basis. As expected, the operator is diagonal with eigenvalues given by the classical values of entropy production Eq. 4. The average is then given by Eq. 7 and holds even for non-full-rank channels, due to the continuity of Umegaki’s relative entropy [46, 47, 48].
An important point to emphasize is that the classical case is recovered by a measure-and-prepare channel, not by the quantum identity channel (which is not measure-and-prepare) or any other unitary channel, even if the unitary operator is just a permutation of the elements of the basis. In fact, as we already noticed in Eq. 25, for a unitary channel (thus including the quantum identity channel), the entropy production is zero. Instead, in the classical formula Eq. 7, even if $D_{\textrm{in}}(p\|\pi)=D_{\textrm{out}}(\varphi p\|\varphi\pi)$, we could still have $\left\langle\sigma\right\rangle=D_{\textrm{out}}(\varphi p\|\tau)\geqslant 0$ in general. The formal reason for the apparent paradox is that, even for the identity channel, $[C_{\operatorname{id}},\mathds{1}\otimes\rho^{T}]\neq 0$: our formulation reveals some deep features that differentiate classical and quantum theories. More precisely, the quantum identity channel does not lead to the definition of a joint probability distribution between input and output: this requires a measure-and-prepare channel, and classically all channels are measure-and-prepare.
IV.9 Two-point measurement scheme
We finish with a discussion of the two-point measurement scheme, mostly for legacy reasons, as it was the first proposal to bring fluctuation theorems to the quantum realm [24]. We describe it in a generalized version introduced in Ref. [11]. Having received a classical label $i$ (in the original, the outcome of the first energy measurement), one prepares the state $\rho_{i}$ (in the original, the post-measurement energy eigenstate). If the classical label is distributed according to the distribution $p_{i}$, the input state is $\rho=\sum_{i}p_{i}\rho_{i}\otimes|i\rangle\!\!\!\;\langle i|$, where we keep track of the label in a dedicated register. The prior $\gamma$ must be of the same form, only with a different distribution $\pi_{i}$. Thus $[\rho,\gamma]=0$.
The state then may undergo a CPTP map (in the original, a unitary evolution), but for notational simplicity we can redefine the $\rho_{i}$ to take that evolution into account. Eventually, a measurement (in the original, the second energy measurement) is performed on the system, leading to $\mathcal{E}(\rho)=\sum_{i}p_{i}\mathcal{M}(\rho_{i})\otimes|i\rangle\!\!\!\;\langle i|=\sum_{i,j}p_{i}\operatorname{Tr}\!\left[\rho_{i}\Pi_{j}\right]|j\rangle\!\!\!\;\langle j|\otimes|i\rangle\!\!\!\;\langle i|$. This is a classical process with $\varphi(j|i)=\operatorname{Tr}\!\left[\rho_{i}\Pi_{j}\right]$. Given the nature of the process, it is natural to assume $\tau=\sum_{j}\tau_{j}|j\rangle\!\!\!\;\langle j|$. Thus, in this context, the two-point measurement scheme and its generalization can both be treated as effectively classical processes.
V Case studies
To gain some intuition about our results, we consider a family of collisional models that describe thermalization [49] or, more generally, “homogenization” [50]. In these models, a qubit undergoes $n$ iterations of the same channel $\mathcal{N}$. The channel is defined by preparing an ancilla qubit in a reference state $\xi$ and making it interact with the system through an interaction that preserves the number of excitations in the eigenbasis of $\xi$, denoted {|$\{\ket{0},\ket{1}\}$. Under these conditions, $\xi$ is the fixed point of the channel, and $\mathcal{N}^{n}(\rho)\rightarrow\xi$ for all $\rho$ in the limit of large $n$. To keep the number of parameters limited, we focus on the most symmetric one among such models, that uses the only interaction that preserves the number of excitations in all bases, namely $H_{\textrm{int}}\propto XX+YY+ZZ$, which generates the unitary
$U=\cos\phi\,\mathbb{I}+i(\sin\phi)\,U_{\textrm{swap}}\;.$
Note that the $n=1$ case is thus the same example studied in Ref. [6], motivated precisely by the fact that the non-commuting “charges” defined by the Pauli matrices are conserved.
The dynamics of the system qubit after $n$ collisions is given by
$\langle 0|\mathcal{N}^{n}(\rho)|0\rangle$ $=c^{2n}\langle 0|\rho|0\rangle+(1-c^{2n})\langle 0|\xi|0\rangle\,,$ (35)
$\langle 0|\mathcal{N}^{n}(\rho)|1\rangle$ $=k^{n}\,\langle 0|\rho|1\rangle$
where $c=\cos\phi$ and with [^3] Compared to Refs. [49, 50], we denote here by $k$ what was written there as $\lambda\cos\phi$.
$k=\frac{1}{2}\Big{[}1+\cos 2\phi+i\sin 2\phi\big{(}2\langle 0|\xi|0\rangle-1\big{)}\Big{]}\;.$
We are going to plot case studies of the average entropy production for the family of channels $\mathcal{N}^{n}$, for varying $n$: they are full-rank channels as long as $\xi$ is not pure, so we can use (IV.5). Our plots represent the $(x,z)$ plane of the Bloch sphere of $\rho$, for given choices of $\gamma$ and $\tau$. The heatmap shows the value of entropy production.
First, in Figure 2 we take a look at the difference between our two definitions Eq. 8 and Section IV.5 of the average entropy production. For the term that depends on the input states, we know that $D_{\textrm{BS}}(\rho\|\gamma)\geqslant D(\rho\|\gamma)$, and so the difference is always non-negative (left panel). Instead, the difference of the terms that depend on the output states, i.e, $\textrm{Tr}\left[\mathcal{N}(\rho)\log\left(\mathcal{N}(\gamma)^{-1/2}\tau\mathcal{N}(\gamma)^{-1/2}\right)\right]-\left[D(\mathcal{N}(\rho)\|\mathcal{N}(\gamma))-D(\mathcal{N}(\rho)\|\tau)\right]$ can have both signs (center panel). The total difference (right panel) can also be positive or negative.
Figure 2: Comparison of our two definitions of average entropy production, Eqs. (8) and (IV.5). Left panel: difference between the terms depending on the input states. Center panel: difference between the terms depending on the output states. Right panel: total difference, i.e., Section IV.5 minus Eq. 8. For the plot, the channel is defined by $\langle 0|\xi|0\rangle=0.95$, $\phi=0.4$, and $n=1$. The Bloch vector of $\gamma$ is $(0.9,0,0)$; that of $\tau$ is $(-2/3,0,-2/3)$.
The next case study, reported in Figure 3, considers the most standard situation: the prior $\gamma$ is taken to be the fixed point $\xi$ of the channel (the Gibbs state, in the thermal interpretation); and the reverse process is made to start from $\tau=\mathcal{N}^{n}(\rho)$. Our prediction is plotted in the top row for three values of $n$; the corresponding value of the more conventional formula $D(\rho\|\xi)-D(\mathcal{N}^{n}(\rho)\|\xi)$, which would follow from Eq. 8, is plotted in the bottom row. As noted in Eq. 26, we generically predict slightly higher values. The two expressions coincide for any $n$ on the $z$ axis ($k=0$): indeed, first note that $\mathcal{N}^{n}(\rho)$ remains diagonal, so $[\tau,\mathcal{N}^{n}(\gamma)]\equiv[\mathcal{N}^{n}(\rho),\xi]=0$, and entropy production is given by (24). In this expression, the last term is zero, and $D_{\textrm{BS}}(\rho\|\gamma)=D(\rho\|\gamma)$, since $[\rho,\gamma]=0$. In the limit $n\rightarrow\infty$ we also have equality in Eq. 26, because then $\tau=\mathcal{N}^{n}(\rho)\rightarrow\xi$. As a global observation, for any initial state $\rho$ we observe that the entropy does not decrease with increasing number of collisions.
Figure 3: Average entropy production for $\gamma=\xi=\mathcal{N}^{n}(\gamma)$ (the dot in the figure) and $\tau=\mathcal{N}^{n}(\rho)$. For the numerics, $\langle 0|\xi|0\rangle=0.9$, $\phi=0.2$, and $n$ given in the figure. Top row: average entropy production as given in Section IV.5. Bottom row: average entropy production as postulated in Eq. 8. The two formulas agree when $[\tau,\mathcal{N}^{n}(\gamma)]=0$ and $[\rho,\gamma]=0$ hold, which is true for all $n$ on the $z$ axis.
For our last case study, shown in Figure 4, we set $\tau=\xi$, with $\gamma$ diagonal in a different basis. After many collisions, $\mathcal{N}^{n}(\rho)\approx\mathcal{N}^{n}(\gamma)\approx\tau$: thus, the second contribution to entropy production (IV.5) tends to zero, and entropy production is dominated by $D_{\textrm{BS}}(\rho\|\gamma)$. In particular, we see that for some $\rho$ around $\gamma$, entropy production decreases as the number of collisions increases. This is not a quantum effect and can happen in the classical case. For example, in Eq. 7, consider $\varphi$ to be a permutation, so that the first two terms cancel: it is quite possible that some iterations of the permutation bring the initial state closer to $\tau$, thus decreasing the entropy over the sequential process.
Figure 4: Average entropy production as defined in Section IV.5, for $\tau=\xi$ and $\gamma$ diagonal in a different basis. For the numerics, $\langle 0|\xi|0\rangle=0.9$, $\phi=0.2$, and $n$ as indicated in the figure. The Bloch vector of $\gamma$ is $(0.8,0,-0.2)$. As the number $n$ of collisions increases, the entropy production tends to $D_{\textrm{BS}}(\rho\|\gamma)$, as clearly shown by the level curves in the right panel, which center around $\gamma$ and become less and less dependent on $\tau$.
VI Conclusion
We have proposed two definitions of fully quantum entropy production, built on the formalism of stochastic thermodynamics, where the subjective element is approached with Bayesian formalism. Definition 1 [Eq. 8] covers all the cases one might want for the average entropy production, and indeed it is the generalization of many versions that have appeared in the literature. It just does not say what is averaged.
The bulk of the paper was devoted to an approach that was not attempted previously: the construction of an entropy production operator based on quantum versions of the input-output processes (Definition 2). The operator $\Sigma[Q_{F},Q_{R}^{\gamma}]$ that we constructed [Eq. 13] is Hermitian, has non-negative average value, and formally fulfills Jarzynski equality for every process. Its average recovers Definition 1 [Eq. 24] when $[\mathcal{E}(\gamma),\tau]=0$ and the channel is full rank, up to the contribution at the input being measured by the Belavkin-Staszewski relative entropy instead of the more common Umegaki one.
We don’t see this work as the last word, but rather as a starting point for the program sketched in the introduction: having a systematic definition of entropy, valid for all processes, that includes the subjective element through a fixed mathematical structure. Let us finish with a few open problems.
First of all, we were not able to find an entropy production operator, whose average is always given by the very pleasant Definition 1. Perhaps it can be found—or perhaps, the fully quantum average entropy production must differ in general. As a hint that the latter may be the case, we bring up another work of ours, the derivation of quantum Bayesian retrodiction from a minimum change principle [37]. The map derived there coincides with the Petz map when $[\mathcal{E}(\gamma),\tau]=0$: intriguingly, this is the same condition under which average entropy production falls back onto the familiar form. We have considered the possibility of defining the reverse quantum process not with the Petz map, but with the general expression found in that work. This would amount at replacing Eq. 12 with
$Q_{R}^{\gamma}(\tau)$ (36)
$=\sqrt{C_{\mathcal{E}}}\,\left(\left[\sqrt{\tau}(\sqrt{\tau}\mathcal{E}(\gamma)\sqrt{\tau})^{-1/2}\sqrt{\tau}\right]^{2}\otimes\gamma^{T}\right)\,\sqrt{C_{\mathcal{E}}}\,.$
Only the explicit formula (IV.5) would have to be modified; all the other properties remain identical, including the formulas obtained when $[\mathcal{E}(\gamma),\tau]=0$.
Another open question relates to the property of “locality in time” that is a signature of Bayesian inference for classical entropy production [12]. In both our definitions, locality in time is preserved for the average; but in the second definition, at the operator level it holds only for classical-quantum channels. One may want an entropy production operator that satisfies this condition for all processes. For instance, in hindsight one may declare the r.h.s. of Section IV.4 to be the entropy production operator: the eigenvalues are the same, but we have not been able to obtain that expression starting from input-output processes.
One may eventually find a definition that ticks all the desiderata. But it is more likely that, in a fully quantum definition of entropy production, we shall have to give up something we find “natural”. Approaches through quasi-probabilities have allowed entropy to take complex values [4, 6]; while our entropy takes real values, the concessions we may have to make have just been listed. Finding out what entropy means in a fully quantum setting would take us to the heart of quantum thermodynamics.
Acknowledgments
This work is supported by the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant; and by the Ministry of Education, Singapore, under the Tier 2 grant “Bayesian approach to irreversibility” (Grant No. MOE-T2EP50123-0002). F. B. acknowledges support from MEXT Quantum Leap Flagship Program (MEXT QLEAP) Grant No. JPMXS0120319794, from MEXT-JSPS Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe” No. 21H05183, and from JSPS KAKENHI, Grants No. 20K03746 and No. 23K03230.
Appendix A Physical interpretation of $Q_{F}$ and $Q_{R}^{\gamma}$
By Stinespring’s dilation theorem [51], any channel can be written as an isometric channel from $A$ to a composite system $BE$ consisting of the output $B$ and an environment system $E$, followed by tracing out system $E$:
$\mathcal{E}(\rho)=\mathrm{Tr}_{E}[V\rho V^{\dagger}]$ (37)
where $V$ is an isometry from ${\mathcal{H}}_{A}$ to ${\mathcal{H}}_{B}\otimes{\mathcal{H}}_{E}$. The complementary channel [52] is obtained by tracing out the original output system $B$, while keeping the environment system $E$:
$\mathcal{E}^{c}(\rho)\equiv\mathrm{Tr}_{B}[V\rho V^{\dagger}]\,.$ (38)
The complementary channel depends on the isometry $V$ used in the dilation, but is unique up to an isometry on the environment system [53, 54].
We first show that $Q_{F}$ is a complementary channel of $\mathcal{E}$. Choose ${\mathcal{H}}_{E}={\mathcal{H}}_{B^{\prime}}\otimes{\mathcal{H}}_{A^{\prime}}$ with ${\mathcal{H}}_{A^{\prime}}\cong{\mathcal{H}}_{A}$ and ${\mathcal{H}}_{B^{\prime}}\cong{\mathcal{H}}_{B}$. Let $\mathds{1}_{A^{\prime}\leftarrow A}$ be the isomorphism from $A$ to $A^{\prime}$ and $\mathds{1}_{A\leftarrow A^{\prime}}$ be its inverse, and $\mathds{1}_{B^{\prime}\leftarrow B}$ be the isomorphism from $B$ to $B^{\prime}$ and $\mathds{1}_{B\leftarrow B^{\prime}}$ be its inverse. In the following, we use $\rho_{A^{\prime}}\equiv\mathds{1}_{A^{\prime}\leftarrow A}\rho\mathds{1}_{A\leftarrow A^{\prime}}$ to denote the operator on system $A^{\prime}$ corresponding to $\rho$. Similar notations are also used for other operators such as $\mathcal{E}(\gamma)_{B^{\prime}}\equiv\mathds{1}_{B^{\prime}\leftarrow B}\mathcal{E}(\gamma)\mathds{1}_{B\leftarrow B^{\prime}}$ and $\tau_{B^{\prime}}\equiv\mathds{1}_{B^{\prime}\leftarrow B}\tau\mathds{1}_{B\leftarrow B^{\prime}}$. Define
$V\equiv(\mathds{1}_{B}\otimes\sqrt{C_{\mathcal{E}}^{T}})(\ket{\Phi^{+}}_{BB^{\prime}}\otimes\mathds{1}_{A^{\prime}\leftarrow A}):{\mathcal{H}}_{A}\to{\mathcal{H}}_{B}\otimes{\mathcal{H}}_{E}$ (39)
[…]
$V^{\dagger}V=(\bra{\Phi^{+}}_{BB^{\prime}}\otimes\mathds{1}_{A\leftarrow A^{\prime}})(\mathds{1}_{B}\otimes{C_{\mathcal{E}}^{T}})(\ket{\Phi^{+}}_{BB^{\prime}}\otimes\mathds{1}_{A^{\prime}\leftarrow A})=\mathds{1}_{A\leftarrow A^{\prime}}\mathrm{Tr}_{B^{\prime}}[C_{\mathcal{E}}]\mathds{1}_{A^{\prime}\leftarrow A}=\mathds{1}_{A}\,.$ (40)
We will see that with this definition, up to isomorphisms between $A$ and $A^{\prime}$ and between $B$ and $B^{\prime}$,
$\mathcal{E}(\rho)$ $=\mathrm{Tr}_{E}[V\rho V^{\dagger}]\,,$ (41)
$Q_{F}^{T}\equiv Q_{F}(\rho)^{T}=\mathcal{E}^{c}(\rho)$ $\equiv\mathrm{Tr}_{B}[V\rho V^{\dagger}]\,.$ (42)
Eq. 41 implies that $V$ is a correct dilation of $\mathcal{E}$, and Eq. 42 shows that $Q_{F}$ can be viewed as a complementary channel of $\mathcal{E}$ up to a transpose.
$\mathrm{Tr}_{E}[V\rho V^{\dagger}]$ $=\mathrm{Tr}_{E}\left[(\mathds{1}_{B}\otimes\sqrt{C_{\mathcal{E}}^{T}})(|\Phi^{+}\rangle\!\!\!\;\langle\Phi^{+}|_{BB^{\prime}}\otimes\rho_{A^{\prime}})(\mathds{1}_{B}\otimes\sqrt{C_{\mathcal{E}}^{T}})\right]$ (43)
$=\mathrm{Tr}_{E}\left[(\mathds{1}_{B}\otimes{C_{\mathcal{E}}^{T}})(|\Phi^{+}\rangle\!\!\!\;\langle\Phi^{+}|_{BB^{\prime}}\otimes\rho_{A^{\prime}})\right]$ (44)
$=\mathrm{Tr}_{B^{\prime}}\left[|\Phi^{+}\rangle\!\!\!\;\langle\Phi^{+}|_{BB^{\prime}}\left(\mathds{1}_{B}\otimes\mathrm{Tr}_{A^{\prime}}\left[{C_{\mathcal{E}}^{T}}(\mathds{1}_{B^{\prime}}\otimes\rho_{A^{\prime}})\right]\right)\right]$ (45)
$=\mathds{1}_{B\leftarrow B^{\prime}}\mathrm{Tr}_{A^{\prime}}\left[{C_{\mathcal{E}}^{T}}(\mathds{1}_{B^{\prime}}\otimes\rho_{A^{\prime}})\right]^{T}\mathds{1}_{B^{\prime}\leftarrow B}$ (46)
$=\mathcal{E}(\rho)$ (47)
In Eq. 46, we have used $\mathrm{Tr}_{B}[|\Phi^{+}\rangle\!\!\!\;\langle\Phi^{+}|_{BB^{\prime}}(\mathds{1}_{B}\otimes X)]=\mathds{1}_{B\leftarrow B^{\prime}}X^{T}\mathds{1}_{B^{\prime}\leftarrow B}$, where the transpose is defined on the basis {|$\{\ket{i}_{B^{\prime}}\}$. Now we compute the marginal state on the environment system.
$\mathrm{Tr}_{B}[V\rho V^{\dagger}]$ $=\mathrm{Tr}_{B}\left[(\mathds{1}_{B}\otimes\sqrt{C_{\mathcal{E}}^{T}})(|\Phi^{+}\rangle\!\!\!\;\langle\Phi^{+}|_{BB^{\prime}}\otimes\rho_{A^{\prime}})(\mathds{1}_{B}\otimes\sqrt{C_{\mathcal{E}}^{T}})\right]$ (48)
$=\sqrt{C_{\mathcal{E}}^{T}}(\mathds{1}_{B^{\prime}}\otimes\rho_{A^{\prime}})\sqrt{C_{\mathcal{E}}^{T}}$ (49)
$=Q_{F}(\rho)^{T}_{B^{\prime}A^{\prime}}$ (50)
This shows Eq. 42.
Next, we show that $Q_{R}$ is a complementary map of the reverse process $\mathcal{R}_{\mathcal{E}}^{\gamma}$. We define the following isometry
$V\equiv(\sqrt{C_{\mathcal{E}}}\otimes\sqrt{\gamma})(\mathds{1}_{B^{\prime}\leftarrow B}\mathcal{E}(\gamma)^{-1/2}\otimes\ket{\Phi^{+}}_{A^{\prime}A}):{\mathcal{H}}_{B}\to{\mathcal{H}}_{E}\otimes{\mathcal{H}}_{A}$ (51)
and will show that
$\mathcal{R}_{\mathcal{E}}^{\gamma}(\tau)$ $=\mathrm{Tr}_{E}[V\tau V^{\dagger}]\,,$ (52)
$Q_{R}^{\gamma}(\tau)=(\mathcal{R}_{\mathcal{E}}^{\gamma})^{c}(\tau)$ $\equiv\mathrm{Tr}_{A}[V\tau V^{\dagger}]\,.$ (53)
First, $V$ is an isometry because
$V^{\dagger}V$ $=(\mathcal{E}(\gamma)^{-1/2}\mathds{1}_{B\leftarrow B^{\prime}}\otimes\bra{\Phi^{+}}_{A^{\prime}A})({C_{\mathcal{E}}}\otimes{\gamma})(\mathds{1}_{B^{\prime}\leftarrow B}\mathcal{E}(\gamma)^{-1/2}\otimes\ket{\Phi^{+}}_{A^{\prime}A})$ (54)
$=\mathcal{E}(\gamma)^{-1/2}\mathds{1}_{B\leftarrow B^{\prime}}\mathrm{Tr}_{A^{\prime}}\left[{C_{\mathcal{E}}}(\mathds{1}_{B^{\prime}}\otimes{\gamma}^{T})\right]\mathds{1}_{B^{\prime}\leftarrow B}\mathcal{E}(\gamma)^{-1/2}$ (55)
$=\mathcal{E}(\gamma)^{-1/2}\mathcal{E}(\gamma)\mathcal{E}(\gamma)^{-1/2}$ (56)
$=\mathds{1}_{B}$ (57)
Second,
$\mathrm{Tr}_{E}[V\tau V^{\dagger}]$ $=\mathrm{Tr}_{E}\left[(\sqrt{C_{\mathcal{E}}}\otimes\sqrt{\gamma})\left(\mathcal{E}(\gamma)^{-1/2}_{B^{\prime}}\tau_{B^{\prime}}\mathcal{E}(\gamma)^{-1/2}_{B^{\prime}}\otimes|\Phi^{+}\rangle\!\!\!\;\langle\Phi^{+}|_{A^{\prime}A}\right)(\sqrt{C_{\mathcal{E}}}\otimes\sqrt{\gamma})\right]$ (58)
$=\sqrt{\gamma}\,\mathrm{Tr}_{E}\left[({C_{\mathcal{E}}}\otimes\mathds{1}_{A})\left(\mathcal{E}(\gamma)^{-1/2}_{B^{\prime}}\tau_{B^{\prime}}\mathcal{E}(\gamma)^{-1/2}_{B^{\prime}}\otimes|\Phi^{+}\rangle\!\!\!\;\langle\Phi^{+}|_{A^{\prime}A}\right)\right]\sqrt{\gamma}$ (59)
$=\sqrt{\gamma}\,\mathrm{Tr}_{A^{\prime}}\left[\left(\mathrm{Tr}_{B^{\prime}}\left[{C_{\mathcal{E}}}\left(\mathcal{E}(\gamma)^{-1/2}_{B^{\prime}}\tau_{B^{\prime}}\mathcal{E}(\gamma)^{-1/2}_{B^{\prime}}\otimes\mathds{1}_{A^{\prime}}\right)\right]\otimes\mathds{1}_{A}\right)\otimes|\Phi^{+}\rangle\!\!\!\;\langle\Phi^{+}|_{A^{\prime}A}\right]\sqrt{\gamma}$ (60)
$=\sqrt{\gamma}\,\mathrm{Tr}_{A^{\prime}}\left[\left(\mathcal{E}^{\dagger}\left[\mathcal{E}(\gamma)^{-1/2}\tau\mathcal{E}(\gamma)^{-1/2}\right]^{T}_{A^{\prime}}\otimes\mathds{1}_{A}\right)\otimes|\Phi^{+}\rangle\!\!\!\;\langle\Phi^{+}|_{A^{\prime}A}\right]\sqrt{\gamma}$ (61)
$=\sqrt{\gamma}\,\mathcal{E}^{\dagger}\left[\mathcal{E}(\gamma)^{-1/2}\tau\mathcal{E}(\gamma)^{-1/2}\right]\sqrt{\gamma}$ (62)
$=\mathcal{R}_{\mathcal{E}}^{\gamma}(\tau)$ (63)
which shows Eq. 52. In Eq. 61, we have used $\mathrm{Tr}_{B}[C_{\mathcal{E}}(\sigma\otimes\mathds{1}_{A})]=\mathcal{E}^{\dagger}(\sigma)^{T}$. Last,
$\mathrm{Tr}_{A}[V\tau V^{\dagger}]$ $=\mathrm{Tr}_{A}\left[(\sqrt{C_{\mathcal{E}}}\otimes\sqrt{\gamma})\left(\mathcal{E}(\gamma)^{-1/2}_{B^{\prime}}\tau_{B^{\prime}}\mathcal{E}(\gamma)^{-1/2}_{B^{\prime}}\otimes|\Phi^{+}\rangle\!\!\!\;\langle\Phi^{+}|_{A^{\prime}A}\right)(\sqrt{C_{\mathcal{E}}}\otimes\sqrt{\gamma})\right]$ (64)
$=\sqrt{C_{\mathcal{E}}}(\mathcal{E}(\gamma)^{-1/2}_{B^{\prime}}\tau_{B^{\prime}}\mathcal{E}(\gamma)^{-1/2}_{B^{\prime}}\otimes\mathds{1}_{A^{\prime}})\,\mathrm{Tr}_{A}\left[(\mathds{1}_{A^{\prime}}\otimes{\gamma})|\Phi^{+}\rangle\!\!\!\;\langle\Phi^{+}|_{A^{\prime}A}\right]\sqrt{C_{\mathcal{E}}}$ (65)
$=\sqrt{C_{\mathcal{E}}}(\mathcal{E}(\gamma)^{-1/2}_{B^{\prime}}\tau_{B^{\prime}}\mathcal{E}(\gamma)^{-1/2}_{B^{\prime}}\otimes\mathds{1}_{A^{\prime}})\gamma^{T}_{A^{\prime}}\sqrt{C_{\mathcal{E}}}$ (66)
$=\sqrt{C_{\mathcal{E}}}(\mathcal{E}(\gamma)^{-1/2}_{B^{\prime}}\tau_{B^{\prime}}\mathcal{E}(\gamma)^{-1/2}_{B^{\prime}}\otimes\gamma^{T}_{A^{\prime}})\sqrt{C_{\mathcal{E}}}$ (67)
$=Q_{R}^{\gamma}(\tau)_{B^{\prime}A^{\prime}}$ (68)
This shows Eq. 53.
Appendix B Proof of Eq. 26
The main tool we use to prove inequalities is the following result in Ref. [55]:
Lemma 1.
For positive operators $X,Y,Z$, with $\mathrm{Tr}[X]=\mathrm{Tr}[Z]$, and all $p>0$,
$\mathrm{Tr}[X\log(Y^{p/2}Z^{p}Y^{p/2})]\leqslant\mathrm{Tr}[X(\log X^{p}+\log Y^{p})]\leqslant\mathrm{Tr}[X\log(X^{p/2}Y^{p}X^{p/2})]$ (69)
The first inequality is strict unless $[Y,Z]=0$, and the second inequality becomes equality if and only if $[X,Y]=0$.
We use this to prove Eq. 26. By Lemma 1 with $X=Z=\mathcal{E}(\rho)$, $Y=\mathcal{E}(\gamma)^{-1}$ and $p=1$,
$\textrm{Tr}\left[\mathcal{E}(\rho)\log\left(\mathcal{E}(\gamma)^{-1/2}\mathcal{E}(\rho)\mathcal{E}(\gamma)^{-1/2}\right)\right]\leqslant\mathrm{Tr}\left[\mathcal{E}(\rho)(\log\mathcal{E}(\rho)-\log\mathcal{E}(\gamma)\right]=D(\mathcal{E}(\rho)\|\mathcal{E}(\gamma))\,.$ (70)
With this inequality, for $\tau=\mathcal{E}(\rho)$,
$\left\langle\Sigma[Q_{F},Q_{R}^{\gamma}]\right\rangle_{F}$ $=D_{\textrm{BS}}(\rho\|\gamma)-\textrm{Tr}\left[\mathcal{E}(\rho)\log\left(\mathcal{E}(\gamma)^{-1/2}\mathcal{E}(\rho)\mathcal{E}(\gamma)^{-1/2}\right)\right]$ (71)
$\geqslant D_{\textrm{BS}}(\rho\|\gamma)-D(\mathcal{E}(\rho)\|\mathcal{E}(\gamma))$ (72)
Appendix C Superadditivity of average entropy production
By Section IV.5,
$\left\langle\Sigma_{1}\right\rangle_{F,1}+\left\langle\Sigma_{2}\right\rangle_{F,2}-\left\langle\Sigma_{12}\right\rangle_{F}$ (73)
$=D_{\textrm{BS}}(\rho\|\gamma)-\textrm{Tr}\left[\mathcal{E}_{1}(\rho)\log\left(\mathcal{E}_{1}(\gamma)^{-1/2}\tau_{1}\mathcal{E}_{1}(\gamma)^{-1/2}\right)\right]$
$\mathrel{\phantom{=}}+D_{\textrm{BS}}(\mathcal{E}_{1}(\rho)\|\mathcal{E}_{1}(\gamma))-\textrm{Tr}\left[\mathcal{E}_{2}[\mathcal{E}_{1}(\rho)]\log\left(\mathcal{E}_{2}[\mathcal{E}_{1}(\gamma)]^{-1/2}\tau_{2}\mathcal{E}_{2}[\mathcal{E}_{1}(\gamma)]^{-1/2}\right)\right]$
$\mathrel{\phantom{=}}-\left(D_{\textrm{BS}}(\rho\|\gamma)-\textrm{Tr}\left[\mathcal{E}_{2}[\mathcal{E}_{1}(\rho)]\log\left(\mathcal{E}_{2}[\mathcal{E}_{1}(\gamma)]^{-1/2}\tau_{2}\mathcal{E}_{2}[\mathcal{E}_{1}(\gamma)]^{-1/2}\right)\right]\right)$ (74)
$=D_{\textrm{BS}}(\mathcal{E}_{1}(\rho)\|\mathcal{E}_{1}(\gamma))-\textrm{Tr}\left[\mathcal{E}_{1}(\rho)\log\left(\mathcal{E}_{1}(\gamma)^{-1/2}\tau_{1}\mathcal{E}_{1}(\gamma)^{-1/2}\right)\right]$ (75)
The last line is no less than zero by taking $X=\mathcal{E}_{1}(\rho)$, $Y=\mathcal{E}_{1}(\gamma)^{-1}$, $Z=\tau_{1}$ and $p=1$ in Lemma 1.
This lemma also indicates that, for Eq. 75 to be zero, one needs $[\mathcal{E}_{1}(\rho),\mathcal{E}_{1}(\gamma)]=[\mathcal{E}_{1}(\gamma),\tau_{1}]=0$, in which condition Eq. 75 becomes $D(\mathcal{E}_{1}(\rho)\|\tau_{1})$, which is zero if and only if $\tau_{1}=\mathcal{E}_{1}(\rho)$. Therefore, $\left\langle\Sigma_{1}\right\rangle_{F,1}+\left\langle\Sigma_{2}\right\rangle_{F,2}=\left\langle\Sigma_{12}\right\rangle_{F}$ if and only if $\tau_{1}=\mathcal{E}_{1}(\rho)$ and $[\mathcal{E}_{1}(\rho),\mathcal{E}_{1}(\gamma)]=0$.

### v2

@@
 To go beyond the mere average, we construct an entropy production operator, that generalizes the value of entropy to the non-commutative case.
-The operator is Hermitian; its average is non-negative, so that the Second Law takes its usual form $\left\langle\Sigma\right\rangle\geqslant 0$; it formally satisfies the Jarzynski equality; and recovers the classical expression for classical processes.
+The operator is Hermitian; its average is non-negative, so that the Second Law takes its usual form $\left\langle\Sigma\right\rangle\geqslant 0$; it formally satisfies the Jarzynski equality and the Crooks fluctuation theorem; and recovers the classical expression for classical processes.
 We also present features whose fully quantum version is not a straightforward translation of the classical one: these may point to the core difference between classical and quantum thermodynamics.
@@
 The average entropy production is then the expectation value of this operator over the physical process.
-Notably, this definition automatically satisfies the Jarzynski equality.
+Notably, this definition satisfies both the Jarzynski equality and Crooks’ fluctuation relation.
 The two definitions are not identical in general, and we illustrate this using a canonical family of channels (Section V).
@@
 These operators, similarly to $\tilde{Q}_{F}$ and $\tilde{Q}_{R}^{\gamma}$, are valid bipartite states, but in addition their dependence on $\rho$ and $\tau$ is linear.
-In fact, as shown in Appendix A, $Q_{F}$ and $Q_{R}^{\gamma}$ are both well-defined quantum channels acting on $\rho$ and on $\tau$, respectively.
+In fact, as shown in Appendix A, $Q_{F}$ (up to transposition) and $Q_{R}^{\gamma}$ are both well-defined quantum channels acting on $\rho$ and on $\tau$, respectively.
 More precisely, $Q_{F}$ is equivalent to the complementary channel of $\mathcal{E}$, up to a transpose.
@@
 whenever $Q_{F}$ and $Q_{R}^{\gamma}$ are supported in the same subspace, with the same formal meaning as in the classical case, namely the normalization of the reverse process.
-By contrast, the generalisation of Crooks’ fluctuation theorem shows a difference between classical and quantum fluctuations in the case of non-commuting processes.
+Next we turn to Crooks fluctuation theorem, a relation between the distributions of values of entropy in the forward and the reverse process.
 Since our operator is Hermitian, we can define the values of entropy production as its eigenvalues:
@@
 $\mathbb{P}_{F}(\Sigma_{k})$ $=\,\langle f_{k}|Q_{F}|f_{k}\rangle\,.$ (17)
-Next we need to define entropy production in the reverse process.
+Next we need to define the entropy production in the reverse process.
 The operator is obviously given by exchanging the roles of $Q_{R}^{\gamma}$ and $Q_{F}$.
-Contrary to the classical case, the identity $\Sigma[Q_{R}^{\gamma},Q_{F}]=-\Sigma[Q_{F},Q_{R}^{\gamma}]$ does not hold in general.
-The eigenvalues are indeed the same with opposite signs, as can be proved notcing that $e^{\Sigma[Q_{F},Q_{R}^{\gamma}]}=\sqrt{Q_{F}}(Q_{R}^{\gamma})^{-1}\sqrt{Q_{F}}$ and $e^{-\Sigma[Q_{R}^{\gamma},Q_{F}]}=(Q_{R}^{\gamma})^{-1/2}Q_{F}(Q_{R}^{\gamma})^{-1/2}$ have the same eigenvalues by polar decomposition.
-The eigenvectors are in general different, and so
+In contrast to the classical case, the operator identity $\Sigma[Q_{R}^{\gamma},Q_{F}]=-\Sigma[Q_{F},Q_{R}^{\gamma}]$ does not hold in general.
+The eigenvalues are the same with opposite signs, but the eigenvectors are in general different (see Appendix B):
 $\Sigma[Q_{R}^{\gamma},Q_{F}]\ket{r_{k}}$ $=\,-\Sigma_{k}\ket{r_{k}}\,.$ (18)
 The statistics of entropy on the reverse process are thus given by
-$\mathbb{P}_{R}(-\Sigma_{k})$ $=\,\langle r_{k}|Q_{R}^{\gamma}|r_{k}\rangle=\,\sum_{j}e^{-\Sigma_{j}}|\langle r_{k}|\sqrt{Q_{F}}|f_{j}\rangle|^{2}\,,$ (19)
-where we have rewritten Eq.
-13 as $Q_{R}^{\gamma}=\sqrt{Q_{F}}e^{-\Sigma[Q_{F},Q_{R}^{\gamma}]}\sqrt{Q_{F}}$.
-Eq.
-19 is the fully quantum extension of the Crooks theorem.
-The classical Crooks expression
-$\mathbb{P}_{R}(-\Sigma_{k})$ $=\,e^{-\Sigma_{k}}\,\mathbb{P}_{F}(\Sigma_{k})$ (20)
-is recovered only when $[Q_{F},Q_{R}^{\gamma}]=0$.
-In the general case, non-commutativity (i.e.
-coherence) impacts the fluctuations of entropy.
-Still, like in the classical case, Jarzynski equality (IV.3) can be derived starting from Eq.
-19, as $\sum_{k}\mathbb{P}_{R}(-\Sigma_{k})=\sum_{j}e^{-\Sigma_{j}}\mathbb{P}_{F}(\Sigma_{j})=1$.
-A comparison between the quantum and classical expressions is shown in Fig.
-1 for the channel that we shall use in Section V.
-One might surmise that quantum fluctuations are always larger than classical ones.
-The figure shows that this is often the case, but not always: for some $\rho$, the standard deviation of the quantum distribution is actually smaller.
-Figure 1: Comparison between the statistics of entropy in the quantum Eq.
-19 and the classical Crooks expression Eq.
-20 for a one-qubit channel (the entropy production operator is of dimension 4).
-The channel $\mathcal{N}$ is defined by Eq.
-35 where $\gamma$ is chosen to be equal to $\xi$, and $\tau=\mathcal{N}(\rho)$.
-For the numerics, we choose $\langle 0|\xi|0\rangle=0.8$, $\phi=0.8$ and $n=1$.
-The plot represents the $(x,z)$ plane of the Bloch sphere of $\rho$.
-The heat map represents the standard deviation of the reverse distribution $\mathbb{P}_{R}$: left, for Eq.
-19 obtained from studying our operator; right, computed from $\mathbb{P}_{F}$ assuming the classical Crooks relation Eq.
-20.
-For most $\rho$, the quantum expression shows a larger standard deviation than the classical guess; but for some, the opposite holds.
+$\mathbb{P}_{R}(-\Sigma_{k})$ $=\,\langle r_{k}|Q_{R}^{\gamma}|r_{k}\rangle\,.$ (19)
+Although it is not immediately obvious, by carefully studying the singular value decomposition of $\sqrt{Q_{F}}(Q_{R}^{\gamma})^{-1/2}$, one can prove (see Appendix B) that the probability distributions just defined in Eqs.
+17 and 19 obey the Crooks relation
+$\mathbb{P}_{R}(-\Sigma_{k})$ $=\,e^{-\Sigma_{k}}\,\mathbb{P}_{F}(\Sigma_{k})\,.$ (20)
 IV.4 Locality in time and the Heisenberg work operator
@@
 $\left\langle\Sigma[Q_{F},Q_{R}^{\gamma}]\right\rangle_{F}\stackrel{{\scriptstyle\tau=\mathcal{E}(\rho)}}{{\geqslant}}D_{\textrm{BS}}(\rho\|\gamma)$ $-D(\mathcal{E}(\rho)\|\mathcal{E}(\gamma))\,.$ (26)
-The proof is in Appendix B.
+The proof is in Appendix C.
 The equality does not hold in general, but it does hold if we further have $[\mathcal{E}(\gamma),\tau]\equiv[\mathcal{E}(\gamma),\mathcal{E}(\rho)]=0$, because then it follows directly from (24).
@@
 Equality holds if and only if $\tau_{1}=\mathcal{E}_{1}(\rho)$ and $[\mathcal{E}_{1}(\rho),\mathcal{E}_{1}(\gamma)]=0$.
-The proof of this property is in Appendix C.
+The proof of this property is in Appendix D.
 It is tempting to interpret the above inequality as a time-like analog of the positivity of mutual information, as Eq. 27 could be rewritten as
@@
 The operator $\Sigma[Q_{F},Q_{R}^{\gamma}]$ that we constructed [Eq.
-13] is Hermitian, has non-negative average value, and formally fulfills Jarzynski equality for every process.
+13] is Hermitian, has non-negative average value, and formally fulfills Jarzynski equality and the Crooks fluctuation theorem for every process.
 Its average recovers Definition 1 [Eq.
@@
 As a hint that the latter may be the case, we bring up another work of ours, the derivation of quantum Bayesian retrodiction from a minimum change principle [37].
-The map derived there coincides with the Petz map when $[\mathcal{E}(\gamma),\tau]=0$: intriguingly, this is the same condition under which average entropy production falls back onto the familiar form.
+The map derived there coincides with the Petz map when $[\mathcal{E}(\gamma),\tau]=0$: intriguingly, this is the same condition under which the average entropy production described here assumes the familiar form.
 We have considered the possibility of defining the reverse quantum process not with the Petz map, but with the general expression found in that work.
@@
 This shows Eq. 53.
-Appendix B Proof of Eq. 26
+Appendix B Proof of quantum Crooks fluctuation theorem
+Here we prove Eq. 20 with the entropy production defined in Eqs. 17 and 19.
+Assume $Q_{F}$ and $Q_{R}^{\gamma}$ to be full rank. Let the singular value decomposition of $\sqrt{Q_{F}}(Q_{R}^{\gamma})^{-1/2}$ be
+$\sqrt{Q_{F}}(Q_{R}^{\gamma})^{-1/2}=USV^{\dagger}\,,$ (69)
+where $U$ and $V$ are unitary operators, and $S$ is a positive definite diagonal operator. They can be written as
+$U=\sum_{k}|f_{k}\rangle\!\!\!\;\langle k|,\quad V=\sum_{k}|r_{k}\rangle\!\!\!\;\langle k|,\quad S=\sum_{k}s_{k}|k\rangle\!\!\!\;\langle k|,\quad s_{k}>0$ (70)
+for some orthonormal basis {|k}$\{\ket{k}\}$.
+{|$\{\ket{f_{k}}\}$ and {|$\{\ket{r_{k}}\}$ are the eigenvectors of $\Sigma[Q_{F},Q_{R}^{\gamma}]$ and $\Sigma[Q_{R}^{\gamma},Q_{F}]$, which can be seen from
+$e^{\Sigma[Q_{F},Q_{R}^{\gamma}]}$ $=\sqrt{Q_{F}}(Q_{R}^{\gamma})^{-1}\sqrt{Q_{F}}=US^{2}U^{\dagger}$ (71)
+$e^{-\Sigma[Q_{R}^{\gamma},Q_{F}]}$ $=(Q_{R}^{\gamma})^{-1/2}Q_{F}(Q_{R}^{\gamma})^{-1/2}=VS^{2}V^{\dagger}$ (72)
+and therefore, $\Sigma[Q_{F},Q_{R}^{\gamma}]=2U(\log S)U^{\dagger}$ and $\Sigma[Q_{R}^{\gamma},Q_{F}]=-2V(\log S)V^{\dagger}$.
+This also verifies that $\Sigma[Q_{F},Q_{R}^{\gamma}]$ and $\Sigma[Q_{R}^{\gamma},Q_{F}]$ have opposite eigenvalues and $\Sigma_{k}=2\log s_{k}$.
+From Eq.
+69 and that $Q_{F},Q_{R}^{\gamma}$ are Hermitian, one gets $Q_{F}=USV^{\dagger}Q_{R}^{\gamma}VSU^{\dagger}$.
+From Eq.
+70, one gets |$\ket{f_{k}}=UV^{\dagger}\ket{r_{k}}$.
+From these, for any $k$,
+$\mathbb{P}_{F}(\Sigma_{k})$ $=\bra{f_{k}}Q_{F}\ket{f_{k}}$ (73)
+$=\bra{r_{k}}VU^{\dagger}\times USV^{\dagger}Q_{R}^{\gamma}VSU^{\dagger}\times UV^{\dagger}\ket{r_{k}}$ (74)
+$=\bra{r_{k}}VSV^{\dagger}Q_{R}^{\gamma}VSV^{\dagger}\ket{r_{k}}$ (75)
+$=s_{k}^{2}\bra{r_{k}}Q_{R}^{\gamma}\ket{r_{k}}$ (76)
+$=s_{k}^{2}~{}\mathbb{P}_{R}(-\Sigma_{k})$ (77)
+noting that |$\ket{r_{k}}$ is the eigenvector of $VSV^{\dagger}$ and $VSV^{\dagger}\ket{r_{k}}=s_{k}\ket{r_{k}}$. Last, since $\Sigma_{k}=2\log s_{k}$, one gets
+$\mathbb{P}_{R}(-\Sigma_{k})=e^{-\Sigma_{k}}\mathbb{P}_{F}(\Sigma_{k})$ (78)
+which is the Crooks fluctuation theorem in Eq. 20.
+Appendix C Proof of Eq. 26
 The main tool we use to prove inequalities is the following result in Ref. [55]:
@@
 $\geqslant D_{\textrm{BS}}(\rho\|\gamma)-D(\mathcal{E}(\rho)\|\mathcal{E}(\gamma))$ (82)
-Appendix C Superadditivity of average entropy production
+Appendix D Superadditivity of average entropy production
 By Section IV.5,